Consider the following production function: q= 4L+K.

Which term describe this production function's returns to scale?

Select one:

a. Decreasing Returns to Scale

b. Increasing Returns to Scale

c. Constant Returns to Scale
Answer #1

Given q=4L+K

Let L and K are increased to xL and xK, then new output is given by

q'=4*(xL)+(xK)=x*(4L+K)=x*q

It shows that if input are increased by a certain proportion, output also get increased by same proportion. It means that the production function exhibits constant returns to scale.

Correct option is

c. Constant Returns to scale.